OpenStudy (anonymous):

Given that the positive numbers p,q,r,s are in G.P. , which of the following must be true? A. kp, kq,kr,ks are in G.P. where k is a non-zero constant. B. a^p , a^q, a^r, a^s are in G.P. where a is a positive constant. C. logp, logq, logr, logs are in A.P.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (caozeyuan):

well, A ofcuz

OpenStudy (anonymous):

why?

OpenStudy (loser66):

what does G.P mean?

OpenStudy (anonymous):

common ratio..

OpenStudy (loser66):

how about AP?

OpenStudy (caozeyuan):

@Loser66 , GP is Geometric Progression or 几何级数 in Chinese, AP is Arithmetic Progression or 算数级数 in CHinese. LOL

OpenStudy (loser66):

I am not chinese, lol

OpenStudy (caozeyuan):

I'm ofcuz

OpenStudy (loser66):

I agree with caozeyuan, it's a

OpenStudy (loser66):

logic is: if p,q,r,s in GP, let d is common ratio, then kp, kq,kr,ks is in another GP whose common ratio is kd

OpenStudy (anonymous):

@caozeyuan oh, you know chinese XD?? @Loser66 I know the definition, but i don't know why B and C are not

OpenStudy (phi):

for choice c: write p, q, r, s as p, kp, k^2 p , k^3 p take the log: log(p), log(kp), log(k^2 p), log(k^3 p) log(p) , log(k) + log(p), log(k^2)+ log(p), log(k^3)+ log(p) or log(p) , log(p) + log(k), log(p) + 2 log(k), log(p) + 3 log(k)

OpenStudy (loser66):

@phi, that means the different is not a constant; contradict to definition of AP, right?

OpenStudy (phi):

for choice c, you get a common difference of log(k), so you have an arithmetic progression

OpenStudy (loser66):

oh, got you

OpenStudy (anonymous):

ohh..

OpenStudy (phi):

for choice b write p, q, r, s as p, kp, k^2 p , k^3 p look at the ratios of two successive terms: a^q / a^p = a^(q-p) = a^(kp-p)= a^(p(k-1)) we need to find the same ratio for a^r/a^q = a^(r-q)= a^(k^2 p - k p) those ratios are different so not a G.P.

OpenStudy (phi):

for A (just to be complete) kq/kp = q/p kr/kq = r/q using write p, q, r, s as p, kp, k^2 p , k^3 p q/p is kp/p = k r/q is k^2 p/ kp = k and so on...

OpenStudy (anonymous):

ohh, clear, that means A and C are correct?